A cab company has computed its mean fare to be $35.00, with a standard deviation of $4.20. Suppose that the cab fares are normally distributed. Complete the following statements. 


(a) Approximately 99.7% of the company's cab rides have fares between $_______ and $_______ 

(b) Approximately _______ of the company's cab rides have fares between $30.80  and $39.20.
Answer #1

a) 99.7% data lies within 3 standard deviations from the mean. Hence,

Interval:

Between $(35 - 3*4.20) and $(35 + 3*4.20)

Between $ 22.40 and $ 47.60

b) z score for 30.80 = (30.80 - 35)/4.20 = -1

z score for 39.20 = (39.20 - 35)/4.20) = 1

So,

Approximate percentage = P(-1 < z < 1) = 68 %